OpenStudy (waheguru):

how to u multiply 2 fractions one being a negative one being a positive

OpenStudy (waheguru):

tell me steps plz

OpenStudy (anonymous):

multiply across and the result will be negative

OpenStudy (anonymous):

here are examples: 1/2*-1/4=-1/8. positive times negative gives negative. -1/2*-1/4=1/8. negative times negative gives positive

OpenStudy (chaise):

You need to either make the denominator the same by multiplying top and bottom of the equation so there is a common denominator. If you are unable to do this you cross multiply. \[a/b+c/d = (ad+bc)/bd\]

OpenStudy (anonymous):

\[\frac{-a}{b}*\frac{c}{d}=\frac{-ac}{bd}\]

OpenStudy (waheguru):

so whenever there is one negative fraction times a positive fractin the answer has 2 be negative?

OpenStudy (anonymous):

2/3 * -4/5 -8/15

OpenStudy (chaise):

Yes, whenever there is one negative multiplied with a positive the answer will ALWAYS be negative.

OpenStudy (waheguru):

thxx soo much